a) with out solving the equation (or factoring) determine the solutions to the equations 7-6x-x^2=0 using only the graph. ( The graph is an upside down U with the right bottom leg at -8.5, the top of the curve is at 15 and the left leg is at 4 .. they intercept at -7,7 and 2)
b. Does this function have a maximum or minumum?
c. What are the coordinates of the vertex in (x,y) form?
d. what is the equation of the axis of symmetry for this parabola?
Breifly explain answers.

Answer by Edwin McCravy(20086)


a) with out solving the equation (or factoring) determine the solutions to the equations 7-6x-x²=0 using only the graph.
The graph of y=7-6x-x² crosses the x-axis at (-7,0) and (1,0) and they are
where y=0, so the solutions of 7-6x-x²=0 are x = -7 and x = 1

b. Does this function have a maximum or minumum?
A maximum because the vertex is the highest point.

c. What are the coordinates of the vertex in (x,y) form?
(-3,16) because the vertex is above -3 on the x-axis and beside 16
on the y-axis.  

d. what is the equation of the axis of symmetry for this parabola?
Every point on the green axis of symmetry is in line with -3 on the
x-axis, so its equation is x = -3.


Edwin